# RUBIES: a complete census of the bright and red distant Universe with JWST/NIRSpec

### v1

@@ A.1 Calibration corrections derived from empty slitlets
We first estimate the 1D sky spectrum of each slitlet by fitting a high-order cubic spline to the intensities of pixels near the expected centers of the open shutters across exposures including that slitlet (typically for GO-2750). Assuming that the shape and normalisation of the sky spectrum should be the same for every slitlet across the detector, these spectra can be used as a highly-multiplexed observation of an (infinitely) extended “standard” source. Using the nominal calibrations from the Calibration Reference Data System (CRDS, jwst_1225.pmap) we find significant variation in the shape of the sky spectra between and within the two NIRCam detectors (Fig. 13, left panels). We derive a shutter- and wavelength-dependent photometric correction of the PRISM spectra by fitting a quadratic 2D polynomial to the spline coefficients of the sky spectra as a function of the MSA shutter row and column indices. This correction reduces the spatial systematics to at where the sky is bright and at where the sky is faintest (Fig. 13, right panels).

### v2

@@ A.1 Calibration corrections derived from empty slitlets
We first estimate the 1D sky spectrum of each slitlet by fitting a high-order cubic spline to the intensities of pixels near the expected centres of the open shutters across exposures including that slitlet (typically for GO-2750). Assuming that the shape and normalisation of the sky spectrum should be the same for every slitlet across the detector, these spectra can be used as a highly-multiplexed observation of an (infinitely) extended “standard” source. Using the nominal calibrations from the Calibration Reference Data System (CRDS, jwst_1225.pmap) we find significant variation in the shape of the sky spectra between and within the two NIRCam detectors (Fig. 13, left panels). We derive a shutter- and wavelength-dependent photometric correction of the PRISM spectra by fitting a quadratic 2D polynomial to the spline coefficients of the sky spectra as a function of the MSA shutter row and column indices. This correction reduces the spatial systematics to at where the sky is bright and at where the sky is faintest (Fig. 13, right panels).